It is necessary to have a 40% antifreeze solution in the radiator of a certain car. The radiator now
has 50 liters of 20% solution. How many liters of this should be drained and replaced with 100%
antifreeze to get the desired strength?

Answer by checkley77(12844)

x+.20(50-x)=.40*50
x=10-.20x=20
.8x=20-10
.8x=10
x=10/.8
x=12.5 liters need to be drained & filled with 100% antifreeze.
Proof:
12.5+.20(50-12.5)=.40*50
12.5+.20*37.5=20
12.5+7.5=20
20=20